• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Benaoum

    Log entry time 00:02:30 on July21,2004

    Entry number 129089

    keyword=Swing Shift Summary



    16:04 changed the IHWP to IN
    16:12 Start run 2999 short run13k
    16:14 Start run 3000 production 80k good data (ended this run DAQ problem).
    17:00 septum Control Stripchart frozen. Rebooted it successfully.
    17:14 start run 3001 production.
    18:14 Asked MCC for Harp Scan.
    18:20 no beam MCC has some problem with FFB.
    18:40 MCC doing Harp Scan. They reduced current to 8uA.
    18:55 MCC finished Harp Scan. See result :
    http://opweb.acc.jlab.org/CSUEApps/elog02/elog.php?logbook=ELOG
    18:57 Asked for 38uA and start run 3002 production 11k (run ended Kent want to put BPMs in forced gain).
    19:06 start run 3003 .
    19:10 MCC has problem to put ON FFB ... They think, it's related to BPMs in forced gain that they just did.
    20:02 end run 3003 with 100k and still FFB on standby.
    20:03 start run 3004. MCC called. They want to put BPMs not in forced gain to have FFB on.
    20:28 Septum Stripchart frozen again. Rebooted septa to gain control.
    20:38 ended run 3004 because FFB is on now. start new run
    20:39 run 3005 production with FFB on.
    21:36 run 3006 production. end run at
    21:34 changed IHWP to OUT.
    21:35 start run 3007 production.
    23:31 start run 3008 still running.......... next shift crew take care of it.